OpenStudy (anonymous):

the first term in a sequence is 400. Every subsequent term is 20 less than half of the immediate preceding term. what is the 4th term in the sequence? A) 10 B) 15 C) 20 D) 25 E) 30

OpenStudy (anonymous):

First term: 400 Second term: 400/2 = 200 -20 = 180 Third term: 180/2 = 90-20 = 70 Fourth term: 70/2 = 35-20 = 15 Choice B :)